# The computational power of discrete chemical reaction networks with bounded executions

### v3

@@ 4 Execution bounded CRDs stably decide all semilinear sets
[…]
{lemmarep}
Every mod set $M=\big{\{}\left(x_{1},\ldots,x_{d}\right)\mid\sum_{i=1}^{d}w_{i}x_{i}\equiv c\bmod m\big{\}}$ is stably decidable by an execution bounded, single-voting CRD with expected stabilization time $\Theta(n\log n)$.
[…]
Proof 4.3.
@@
$\mathcal{D}$ stabilizes in $\Theta(n\log n)$ time: We start with #$\#L=1,\#X=n$ in volume $n$. $n$ reactions must occur before $\mathcal{D}$ terminates. For the first reaction, we have a rate of $\lambda=\frac{n\cdot 1}{n}$, for the last (with only the leader and one $X$ present), our rate will be $\lambda=\frac{1\cdot 1}{n}$. Thus, the expected time for all $n$ reactions to complete is
@@
{lemmarep}
Every threshold set $T=\big{\{}(x_{1},\ldots,x_{d})\mid\sum_{i=1}^{d}w_{i}x_{i}\geq t\big{\}}$ is stably decidable by an execution bounded, single-voting CRD with expected stabilization time $\Theta(n\log n)$.
@@
[…]
Proof 4.4.
@@
$\mathcal{D}$ stabilizes in $\Theta(n\log n)$ time: First, all input species will be converted to $w_{i}$ instances of $P$ or $N$. We run these reactions until no $X_{i}$. As they are independent of molecules other than the reactant, these reactions have a rate of $\lambda=i$, so the expected time until the next reaction is $\frac{1}{i}$. The total time for reactions (5) to complete is therefore $\sum_{i=1}^{n}\frac{1}{i}=\mathit{\Theta}(\log n)$. The time for reactions (6) and (7) on the other hand is asymptotically dominated by the last reaction, where #$\#L=1$ and #$\#B=1$, where $B\in\{P,N\}$, so $\lambda=\frac{1\cdot 1}{n}$. Let $n_{P},n_{N}$ be the counts of $P,N$ and assume without loss of generality $n_{P}\geq n_{N}$. We get:
@@
Reaction (8) will complete in $O(\log n)$ time and is clearly execution bounded since the input $X_{i}$ is finite and not produced in any reaction. Consequently, two separate CRNs run in $\mathit{\Theta}(n\log n)$ time as shown in Section 4 and Section 4. After stabilization of the parallel CRNs, we expect reaction (9) and (10) to happen exactly once. Each molecule involved is a leader and has count $1$ in volume $n$. This leads to a rate of $\lambda=\frac{1\cdot 1}{n}$, so the expected time for one reaction to happen is $O(n)$. It is important to note that reactions (9) and (10) do not result in unbounded executions due to the unanimous vote in parallel CRDs. In both mod sets and threshold sets, the leader changes its vote a maximum of $|\mathbf{i}|$ times, with only ever one leader present at any time. Again, we start with only one $V_{bb}$ voter present initially and no reaction changes the count of voters, making our construction single-voting.
Since semilinear predicates are exactly Boolean combinations of threshold and mod predicates, Sections 4, 4 and 4.5 imply Theorem 4.1.
@@
{thmrep}
Every semilinear set is stably decidable by an execution bounded, all-voting CRD, with expected stabilization time $O(n\log n).$
Proof 4.7.
By Theorem 4.1, every semilinear set is stably decided by a single-voting CRD. We convert this to an all-voting CRD, where every species is required to vote yes or no, by “propagating” the final vote (recorded in the single voter $V^{0}$ voting no or $V^{1}$ voting yes) back to all other molecules. A superscript indicates the “global” decision. The execution boundedness proven in Lemma 4.5 ensures that the leader propagates the final vote only a finite amount of times. For each vote $b\in\{0,1\}$ and each voter $V^{b}$ voting $b$, and all other species $S\in\Lambda\backslash\{V\}$, replace species $S$ with two versions $S^{0}$ and $S^{1}$, and add reactions:
@@ 5 Execution bounded CRCs stably compute all semilinear functions
Similar to [5, 7], we compute semilinear functions by decomposing them into “affine pieces”, which we will show can be computed by execution bounded CRNs and combined by using semilinear predicates to decide which linear function to apply for a given input.333 While this proof generalizes to multivariate output functions as in [5, 7], to simplify notation we focus on single output functions. Multi-valued functions $f:\mathbb{N}^{d}\to\mathbb{N}^{l}$ can be equivalently thought of as $l$ separate single output functions $f_{i}:\mathbb{N}^{d}\to\mathbb{N}$, which can be computed in parallel by independent CRCs.
[…]
@@
[…]
submission,final $f(\mathbf{x})=b+\sum_{i=1}^{k}\mathbf{a}(i)(\mathbf{x}(i)-\mathbf{c}(i))\iff f(\mathbf{x})=b+\frac{1}{d}\sum_{i=1}^{k}n_{i}(\mathbf{x}(i)-\mathbf{c}(i)).$
We say that a partial function $\hat{f}:\mathbb{N}^{k}\rightarrow\mathbb{N}^{2}$ is a diff-representation of $f$ if $\operatorname{dom}f=\operatorname{dom}\hat{f}$ and, for all $\mathbf{x}\in\operatorname{dom}f$, if $\left(y_{P},y_{C}\right)=\hat{f}(\mathbf{x})$, then $f(\mathbf{x})=y_{P}-y_{C}$, and $y_{P}=O(f(\mathbf{x}))$. In other words, $\hat{f}$ represents $f$ as the difference of its two outputs $y_{P}$ and $y_{C}$, with the larger output $y_{P}$ possibly being larger than the original function’s output, but at most a multiplicative constant larger [7].
@@
We require the following result due to Chen, Doty, Soloveichik [5], guaranteeing that any semilinear function can be built from affine partial functions.
Lemma 5.3 ([5]).
@@
[…]
{lemmarep}
Let $f:\mathbb{N}^{d}\rightarrow\mathbb{N}$ be a semilinear function. Then there is a finite set $\big{\{}f_{1}:\mathbb{N}^{d}\rightarrow\mathbb{N},\ldots,f_{m}:\mathbb{N}^{d}\rightarrow\mathbb{N}\big{\}}$ of affine partial functions, where each $\operatorname{dom}f_{i}$ is a linear set, and $\operatorname{dom}f_{i}\cap\operatorname{dom}f_{j}=\emptyset$ for all $i\neq j$, such that, for each $\mathbf{x}\in\mathbb{N}^{d}$, if $f_{i}(\mathbf{x})$ is defined, then $f(\mathbf{x})=f_{i}(\mathbf{x})$, and $\bigcup_{i=1}^{m}\operatorname{dom}f_{i}=\mathbb{N}^{d}$.
Proof 5.4.
By [12, Theorem 2], every semilinear set is a finite union of disjoint fundamental linear sets. The author defines a linear set $L=\big{\{}\mathbf{b}+n_{1}\mathbf{u}_{1}+\ldots+n_{p}\mathbf{u}_{p}\mid n_{1},\ldots,n_{p}\in\mathbb{N}\big{\}}$ as fundamental, if $\mathbf{u}_{1},\dots\mathbf{u}_{p}\in\mathbb{N}^{k}$ span a $p$-dimensional vector space in $\mathbb{R}^{k}$, i.e. all vectors are linearly independent in $\mathbb{R}^{k}$.444 This distinction is significant because not all integer-valued linear sets can be represented using solely linearly independent vectors. An illustrative example is $\mathbf{b}=\mathbf{0},\mathbf{u}_{1}=(1,1,1),\mathbf{u}_{2}=(2,0,1),\mathbf{u}_{3}=(0,2,1)$, as discussed in [5]. The vectors $\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3}$ are not linearly independent in $\mathbb{R}^{3}$, yet this set cannot be expressed with less than three basis vectors. The proof of Lemma 5.3 in [5] shows that each linear set $L_{i}$ comprising the semilinear graph of $f$ corresponds to one partial affine function $f_{i}$. The fact that [12, Theorem 2] lets us assume each $L_{i}$ is disjoint from the others immediately implies that each $\mathrm{dom}\ f_{i}$ is disjoint from the others.
@@
Proof 5.6.
We employ the same construction of [5] with minor alterations. A CRC with input species $\Sigma=\{X_{1},\ldots,X_{d}\}$ and output species $\Gamma=\{Y\}$. By Section 5, we decompose our semilinear function into partial affine functions (with linear, disjoint domains), which can be computed in parallel by Lemma 5.1. Further, we decide which function to use by computing the predicate $\phi_{i}=$$\mathrm{x}\in\operatorname{dom}f_{i}$” (Theorem 4.1). We interpret each $\widehat{Y}_{i}^{P}$ and $\widehat{Y}_{i}^{C}$ as an “inactive” version of “active” output species $Y_{i}^{P}$ and $Y_{i}^{C}$. Let $L_{i}^{Y},L_{i}^{N}$ be the yes and no voters respectively voting whether $\mathbf{x}$ lies in the domain of $i$-th partial function. Now, we convert the function result of the applicable partial affine function to the global output by adding the following reactions for each $i\in\{1,\ldots,m\}$.
@@ 6 Limitations of execution bounded CRNs
The main positive results of the paper (Theorems 4.1 and 5.5) rely on the assumption that valid initial configurations have a single leader (in particular, they are execution bounded only from configurations with a single leader, but not from arbitrary configurations). Section 4 shows that we may assume the CRD deciding a semilinear set is all-voting. However, for the “constructive” results LABEL:{lem:boolean-closure-crd} and 5.5, which compose the output of a CRD $\mathcal{D}$ with downstream computation, using $\mathcal{D}$ as a “subroutine” to stably compute a more complex set/function, the constructions crucially use the assumption that $\mathcal{D}$ is single-voting (i.e., only the leader of $\mathcal{D}$ votes) to argue the resulting composed CRN is execution bounded. In this section we show these assumptions are necessary, proving that execution bounded CRNs without those constraints are severely limited in their computational abilities.
[…]
@@ 6.1 Linear potential functions
[…]
@@
[…]
submission,final $(\forall i)\sum_{S\in\Lambda}[\mathbf{p}_{i}(S)-\mathbf{r}_{i}(S)]v_{S}<0.$ For example, for the reactions $A+A\mathop{\rightarrow}\limits B+C$ and $B+B\mathop{\rightarrow}\limits A$, for each reaction to strictly decrease the potential function $\Phi(\mathbf{x})=v_{A}\mathbf{x}(A)+v_{B}\mathbf{x}(B)+v_{C}\mathbf{x}(C)$, $\Phi$ must satisfy $2v_{A}>v_{B}+v_{C}$ and $2v_{B}>v_{A}$. In this case, $v_{A}=1,v_{B}=1,v_{C}=0$ works.
@@
[…]
full,submission
The following is a variant of Farkas’ Lemma [8], one of several similar “Theorems of the Alternative” stating that exactly one of two different linear systems has a solution. (See [14, Chapter 4] for a list of such theorems.) A proof can be found in [9, Theorem 2.10].
@@
[…]
{corollaryrep}
Let $\mathbf{M}$ be a rational matrix. Exactly one of the following statements is true.
@@
Proof 6.4.
For convenience when we use Section 6.1 in proving Theorem 6.5, we swapped the roles of $\mathbf{u}$ and $\mathbf{v}$ in left- vs. right-multiplication with $\mathbf{M}$; the real-valued version of the statement of Section 6.1 is equivalent to Theorem 6.3 by taking the transpose of $\mathbf{M}$.
@@
Figure 1: Geometric intuition of Section 6.1. A matrix $\mathbf{M}$ has column vectors $\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3}$. The cone of $\mathbf{M}$ is the nonnegative span of these vectors, shown as a faded gray region. Exactly one of two scenarios occurs: a) The cone of $\mathbf{M}$ intersects the first quadrant (nonnegative orthant in higher dimensions) away from the origin, i.e., some semipositive point ($\mathbf{M}\mathbf{u}\geq\mathbf{0}$ above) is a nonnegative linear combination of vectors $\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3}$. b) The cone of $\mathbf{M}$ does not intersect the first quadrant except at the origin. In this case we can draw a dashed line (hyperplane in higher dimensions) separating the cone of $\mathbf{M}$ from the first quadrant. The orthogonal vector $\mathbf{v}\geq\mathbf{0}$ to this line lies in the first quadrant, but $\mathbf{v}\mathbf{M}<0$ means each vector $\mathbf{x}_{i}$ has negative dot product $\mathbf{v}\cdot\mathbf{x}_{i}<0$, i.e., all $\mathbf{x}_{i}$ vectors lie on the other side of the line.
{toappendix}
Although we do not need the following fact, it is worthwhile to observe that, if $\mathbf{M}$ is integer-valued (as in our application), then the solution $\mathbf{u}$ or $\mathbf{v}$ (whichever exists) in Section 6.1 has entries that are at most exponential in $\|\mathbf{M}\|$ i.e., at most exponential in the sum of absolute values of entries of $\mathbf{M}$ (see e.g., [15]). So in particular when we consider $\mathbf{M}$ having small $O(1)$ size entries, this means the solution $\mathbf{u}$ or $\mathbf{v}$ has entries that are at most exponential in the number of rows and columns of $\mathbf{M}$. When $\mathbf{M}$ is a stoichiometric matrix, this corresponds to the number of species and reactions, respectively, of the CRN.
Section 6.1 will help us prove the following theorem characterizing CRNs with bounded executions from all configurations. Theorem 6.5 is used in this paper only to prove Theorems 6.9 and 6.3, but it may also be of independent interest, since it equates a “global, infinitary, difficult-to-check” condition (bounded executions from all configurations) with a “local, easy-to-check” condition (having a linear potential function).
[…]
finalThe following theorem due to Czerner, Guttenberg, Helfrich, and Esparza, is crucial to proving limitations on execution bounded CRNs such as Theorems 6.9 and 6.3.
Theorem 6.5 ([6]).
A CRN has a linear potential function if and only if it is entirely execution bounded.
[…]
full,submission
Proof 6.6.
Let $\mathcal{C}=(\Lambda,R)$ be a CRN. The forward direction is easy: assuming $\mathcal{C}$ has potential function $\Phi$, since each reaction decreases $\Phi$ by at least 1 (see Remark 6.2), from any configuration $\mathbf{x}$, $\mathcal{C}$ can execute at most $\Phi(\mathbf{x})$ reactions while keeping $\Phi$ nonnegative. Thus $\mathcal{C}$ is entirely execution bounded.
To see the reverse direction, assume that $\mathcal{C}$ is execution bounded from every configuration, and let $\mathbf{M}$ be the stoichiometric matrix of $\mathcal{C}$. We claim there is no integer vector $\mathbf{u}\geq\mathbf{0}$ satisfying $\mathbf{M}\mathbf{u}\geqq\mathbf{0}$; for the sake of contradiction suppose otherwise. Interpreting $\mathbf{u}$ as counts of reactions to execute, for any sufficiently large configuration $\mathbf{x}$, all reactions in $\mathbf{u}$ can be applied (in arbitrary order), and the vector $\mathbf{M}\mathbf{u}$ describes the resulting change in species counts, reaching to configuration $\mathbf{y}=\mathbf{x}+\mathbf{M}\mathbf{u}$. Since $\mathbf{M}\mathbf{u}\geqq\mathbf{0}$, this path is self-covering, i.e., $\mathbf{y}\geqq\mathbf{x}$. But since $\mathcal{C}$ is execution bounded from every configuration, by Section 3, $\mathcal{C}$ has no self-covering path from any configuration, a contradiction. This establishes the claim that $\mathbf{M}\mathbf{u}\geqq\mathbf{0}$ has no integer solution $\mathbf{u}\geq\mathbf{0}$.
By Section 6.1, there is an integer vector $\mathbf{v}\geqq\mathbf{0}$ such that $\mathbf{v}\mathbf{M}<\mathbf{0}$. Let $\mathbf{v}\in\mathbb{N}^{\Lambda}$ be the coefficients of a linear function $\Phi:\mathbb{N}^{\Lambda}\to\mathbb{N}$, i.e., $\Phi(\mathbf{x})=\mathbf{v}\cdot\mathbf{x}$. Then the vector $\mathbf{v}\mathbf{M}\in\mathbb{Z}^{R}$ represents the amount $\Phi$ changes by one unit of each reaction, i.e., $\mathbf{v}\mathbf{M}(\alpha)$ is the amount $\Phi$ increases when executing reaction $\alpha$ once. Since $\mathbf{v}\mathbf{M}<\mathbf{0}$, this means that every reaction strictly decreases $\Phi$, i.e., $\Phi$ is a linear potential function for $\mathcal{C}$.
{toappendix}
Remark 6.7.
By employing the real-valued version of Section 6.1, the above proof also shows that Theorem 6.5 holds for the continuous model of CRNs [4], in which species amounts are modeled as continuous nonnegative real concentrations. In this case, a continuous CRN would be defined to be execution bounded from configuration $\mathbf{x}$ if each reaction can be executed by at most a finite (real-valued) amount from $\mathbf{x}$.

### v4

@@ 4 Execution bounded CRDs stably decide all semilinear sets
Lemma 4.3.
lem:mod_sets_decidable Every mod set $M=\big\{\left(x_{1},\ldots,x_{d}\right)\mid\sum_{i=1}^{d}w_{i}x_{i}\equiv c\bmod m\big\}$ is stably decidable by an execution bounded, single-voting CRD with expected stabilization time $\Theta(n\log n)$.
We design a CRD $\mathcal{D}$ with exactly one leader present at all times, cycling through $m$ “states” while consuming the input and accepting on state $c$. Let $\Sigma=\{X_{1},\dots,X_{d}\}$ be the set of input species and start with only one $L_{0}$ leader, i.e. set the initial context $\mathbf{s}(L_{0})=1$ and $\mathbf{s}(S)=0$ for all other species. For each $i\in\{1,\dots,d\},j\in\{0,\dots,m-1\}$ add the following reaction: $X_{i}+L_{j}\rightarrow L_{j+w_{i}\bmod m}.$ Let only $L_{c}$ vote yes and all other species no, i.e. $\Upsilon=\{L_{c}\}$. For any valid initial configuration, $\mathcal{D}$ reaches a stable configuration which votes yes if and only if the input is in the mod set, and no otherwise.
Proof 4.4.
@@
$\mathcal{D}$ stabilizes in $\Theta(n\log n)$ time: We start with #L=1,#X=n\#L=1,\#X=n in volume $n$. $n$ reactions must occur before $\mathcal{D}$ terminates. For the first reaction, we have a rate of $\lambda=\frac{n\cdot 1}{n}$, for the last (with only the leader and one $X$ present), our rate will be $\lambda=\frac{1\cdot 1}{n}$. Thus, the expected time for all $n$ reactions to complete is
@@
Lemma 4.5.
lem:threshold_sets_decidable Every threshold set $T=\big\{(x_{1},\ldots,x_{d})\mid\sum_{i=1}^{d}w_{i}x_{i}\geq t\big\}$ is stably decidable by an execution bounded, single-voting CRD with expected stabilization time $\Theta(n\log n)$.
@@
Finally, initialize the CRD with one $L_{Y}$ and the threshold number $t$ copies of $P$ (or tN-tN if $t$ is negative), i.e. $\mathbf{s}(L_{Y})=1$, $\mathbf{s}(P)=t$ if $t>0$, or $\mathbf{s}(N)=-t$ if $t<0$, and $\mathbf{s}(S)=0$ for all other species. For any valid initial configuration, $\mathcal{D}$ reaches a stable configuration which votes yes if and only if the weighted sum of inputs is above the threshold, and no otherwise.
Proof 4.6.
@@
$\mathcal{D}$ stabilizes in $\Theta(n\log n)$ time: First, all input species will be converted to $w_{i}$ instances of $P$ or $N$. We run these reactions until no $X_{i}$. As they are independent of molecules other than the reactant, these reactions have a rate of $\lambda=i$, so the expected time until the next reaction is $\frac{1}{i}$. The total time for reactions (5) to complete is therefore $\sum_{i=1}^{n}\frac{1}{i}=\mathit{\Theta}(\log n)$. The time for reactions (6) and (7) on the other hand is asymptotically dominated by the last reaction, where #L=1\#L=1 and #B=1\#B=1, where $B\in\{P,N\}$, so $\lambda=\frac{1\cdot 1}{n}$. Let $n_{P},n_{N}$ be the counts of $P,N$ and assume without loss of generality $n_{P}\geq n_{N}$. We get:
@@
Reaction (8) will complete in $O(\log n)$ time and is clearly execution bounded since the input $X_{i}$ is finite and not produced in any reaction. Consequently, two separate CRNs run in $\mathit{\Theta}(n\log n)$ time as shown in LABEL:lem:mod_sets_decidable and LABEL:lem:threshold_sets_decidable. After stabilization of the parallel CRNs, we expect reaction (9) and (10) to happen exactly once. Each molecule involved is a leader and has count $1$ in volume $n$. This leads to a rate of $\lambda=\frac{1\cdot 1}{n}$, so the expected time for one reaction to happen is $O(n)$. It is important to note that reactions (9) and (10) do not result in unbounded executions due to the unanimous vote in parallel CRDs. In both mod sets and threshold sets, the leader changes its vote a maximum of $|\mathbf{i}|$ times, with only ever one leader present at any time. Again, we start with only one $V_{bb}$ voter present initially and no reaction changes the count of voters, making our construction single-voting.
Since semilinear predicates are exactly Boolean combinations of threshold and mod predicates, LABEL:lem:mod_sets_decidable, LABEL:lem:threshold_sets_decidable and 4.7 imply Theorem˜4.1.
We can also prove the same result for all-voting CRDs. Note, however, that such CRDs cannot be “composed” using the constructions of Lemmas˜4.7 and 5.6, which crucially relied on the assumption that the CRDs being used as “subroutines” are single-voting.
Theorem 0.
lem:boolean-closure-crd-all-voting Every semilinear set is stably decidable by an execution bounded, all-voting CRD, with expected stabilization time $O(n\log n).$
Proof 4.10.
@@ 5 Execution bounded CRCs stably compute all semilinear functions
Similar to [Chen2012DeterministicFunction, Doty2013Leaderless], we compute semilinear functions by decomposing them into “affine pieces”, which we will show can be computed by execution bounded CRNs and combined by using semilinear predicates to decide which linear function to apply for a given input. [^2] While this proof generalizes to multivariate output functions as in [Chen2012DeterministicFunction, Doty2013Leaderless], to simplify notation we focus on single output functions. Multi-valued functions $f:\mathbb{N}^{d}\to\mathbb{N}^{l}$ can be equivalently thought of as $l$ separate single output functions $f_{i}:\mathbb{N}^{d}\to\mathbb{N}$, which can be computed in parallel by independent CRCs.
We say a partial function $f:\mathbb{N}^{k}\dashrightarrow\mathbb{N}$ is affine if there exist a vectors $\mathbf{a}\in\mathbb{Q}^{k}$, $\mathbf{c}\in\mathbb{N}^{k}$ with $\mathbf{x}-\mathbf{c}\geq\mathbf{0}$ and nonnegative integer $b\in\mathbb{N}$ such that $f(\mathbf{x})=\mathbf{a}^{\top}(\mathbf{x}-\mathbf{c})+b.$ For a partial function $f$ we write $\operatorname{dom}f$ for the domain of $f$, the set of inputs for which $f$ is defined. This definition of affine function may appear contrived, but the main utility of the definition is that it satisfies LABEL:lem:partial_affine_function_decomposition_disjoint_domains. For convenience, we can ensure to only work with integer valued molecule counts by multiplying by $\frac{1}{d}$ after the dot product, where $d$ may be taken to be the least common multiple of the denominators of the rational coefficients in the original definition such that $n_{i}=d\cdot\mathbf{a}(i)$:
@@
We say that a partial function $\hat{f}:\mathbb{N}^{k}\rightarrow\mathbb{N}^{2}$ is a diff-representation of $f$ if $\operatorname{dom}f=\operatorname{dom}\hat{f}$ and, for all $\mathbf{x}\in\operatorname{dom}f$, if $\left(y_{P},y_{C}\right)=\hat{f}(\mathbf{x})$, then $f(\mathbf{x})=y_{P}-y_{C}$, and $y_{P}=O(f(\mathbf{x}))$. In other words, $\hat{f}$ represents $f$ as the difference of its two outputs $y_{P}$ and $y_{C}$, with the larger output $y_{P}$ possibly being larger than the original function’s output, but at most a multiplicative constant larger [Doty2013Leaderless].
@@
We require the following result due to Chen, Doty, Soloveichik [Chen2012DeterministicFunction], guaranteeing that any semilinear function can be built from affine partial functions.
Lemma 5.3 ([Chen2012DeterministicFunction]).
@@
We strengthen Lemma˜5.3 to show we may assume each $\mathrm{dom}\ f_{i}$ is disjoint from the others. This is needed not only to prove Theorem˜5.6, but to correct the proof of Lemma 4.4 in [Chen2012DeterministicFunction], which implicitly assumed the domains are disjoint.
Lemma 5.4.
lem:partial_affine_function_decomposition_disjoint_domains Let $f:\mathbb{N}^{d}\rightarrow\mathbb{N}$ be a semilinear function. Then there is a finite set $\big\{f_{1}:\mathbb{N}^{d}\rightarrow\mathbb{N},\ldots,f_{m}:\mathbb{N}^{d}\rightarrow\mathbb{N}\big\}$ of affine partial functions, where each $\operatorname{dom}f_{i}$ is a linear set, and $\operatorname{dom}f_{i}\cap\operatorname{dom}f_{j}=\emptyset$ for all $i\neq j$, such that, for each $\mathbf{x}\in\mathbb{N}^{d}$, if $f_{i}(\mathbf{x})$ is defined, then $f(\mathbf{x})=f_{i}(\mathbf{x})$, and $\bigcup_{i=1}^{m}\operatorname{dom}f_{i}=\mathbb{N}^{d}$.
Proof 5.5.
By [Ito1969SemilinearSetsFiniteUnionDisjointLinearSets, Theorem 2], every semilinear set is a finite union of disjoint fundamental linear sets. The author defines a linear set $L=\big\{\mathbf{b}+n_{1}\mathbf{u}_{1}+\ldots+n_{p}\mathbf{u}_{p}\mid n_{1},\ldots,n_{p}\in\mathbb{N}\big\}$ as fundamental, if $\mathbf{u}_{1},\dots\mathbf{u}_{p}\in\mathbb{N}^{k}$ span a $p$-dimensional vector space in $\mathbb{R}^{k}$, i.e. all vectors are linearly independent in $\mathbb{R}^{k}$. [^3] This distinction is significant because not all integer-valued linear sets can be represented using solely linearly independent vectors. An illustrative example is $\mathbf{b}=\mathbf{0},\mathbf{u}_{1}=(1,1,1),\mathbf{u}_{2}=(2,0,1),\mathbf{u}_{3}=(0,2,1)$, as discussed in [Chen2012DeterministicFunction]. The vectors $\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3}$ are not linearly independent in $\mathbb{R}^{3}$, yet this set cannot be expressed with less than three basis vectors. The proof of Lemma˜5.3 in [Chen2012DeterministicFunction] shows that each linear set $L_{i}$ comprising the semilinear graph of $f$ corresponds to one partial affine function $f_{i}$. The fact that [Ito1969SemilinearSetsFiniteUnionDisjointLinearSets, Theorem 2] lets us assume each $L_{i}$ is disjoint from the others immediately implies that each $\mathrm{dom}\ f_{i}$ is disjoint from the others.
@@
Proof 5.7.
We employ the same construction of [Chen2012DeterministicFunction] with minor alterations. A CRC with input species $\Sigma=\{X_{1},\ldots,X_{d}\}$ and output species $\Gamma=\{Y\}$. By LABEL:lem:partial_affine_function_decomposition_disjoint_domains, we decompose our semilinear function into partial affine functions (with linear, disjoint domains), which can be computed in parallel by Lemma˜5.1. Further, we decide which function to use by computing the predicate $\phi_{i}=$$\mathrm{x}\in\operatorname{dom}f_{i}$” (Theorem˜4.1). We interpret each $\widehat{Y}_{i}^{P}$ and $\widehat{Y}_{i}^{C}$ as an “inactive” version of “active” output species $Y_{i}^{P}$ and $Y_{i}^{C}$. Let $L_{i}^{Y},L_{i}^{N}$ be the yes and no voters respectively voting whether $\mathbf{x}$ lies in the domain of $i$-th partial function. Now, we convert the function result of the applicable partial affine function to the global output by adding the following reactions for each $i\in\{1,\ldots,m\}$.
@@ 6 Limitations of execution bounded CRNs
The main positive results of the paper (Theorems˜4.1 and 5.6) rely on the assumption that valid initial configurations have a single leader (in particular, they are execution bounded only from configurations with a single leader, but not from arbitrary configurations). LABEL:lem:boolean-closure-crd-all-voting shows that we may assume the CRD deciding a semilinear set is all-voting. However, for the “constructive” results LABEL:{lem:boolean-closure-crd} and 5.6, which compose the output of a CRD $\mathcal{D}$ with downstream computation, using $\mathcal{D}$ as a “subroutine” to stably compute a more complex set/function, the constructions crucially use the assumption that $\mathcal{D}$ is single-voting (i.e., only the leader of $\mathcal{D}$ votes) to argue the resulting composed CRN is execution bounded. In this section we show these assumptions are necessary, proving that execution bounded CRNs without those constraints are severely limited in their computational abilities.
We show that entirely execution bounded CRNs (from every configuration) can be characterized by a simpler property of having a “linear potential function” that essentially measures how close the CRN is to reaching a terminal configuration. This result was essentially proven independently by Czerner, Guttenberg, Helfrich, and Esparza [czerner2024fast] in the model of population protocols. We give a self-contained proof since we technically need the more general CRN model, although our proof appears to follow their same technique of using a Farkas-like lemma. We use this characterization to prove that entirely execution bounded CRNs can stably decide only limited semilinear predicates (eventually constant, Definition˜6.9), assuming all species vote, and that molecular counts cannot decrease to $O(1)$ in stable configurations (see Definition˜6.6).
@@ 6.1 Linear potential functions
A CRN may or may not have a linear potential function. Although it is not straightforward to “syntactically check” a CRN to see if has a linear potential function, it is efficiently decidable: a CRN has a linear potential function if and only if the following system of linear inequalities has a solution (which can be solved in polynomial time using linear programming techniques; the variables to solve for are the $v_{S}$ for each $S\in\Lambda$), where the $i$’th reaction has reactants $\mathbf{r}_{i}$ and products $\mathbf{p}_{i}$, and species $S\in\Lambda$ has mass $v_{S}\geq 0$:
@@
For example, for the reactions $A+A\mathop{\rightarrow}\limits B+C$ and $B+B\mathop{\rightarrow}\limits A$, for each reaction to strictly decrease the potential function $\Phi(\mathbf{x})=v_{A}\mathbf{x}(A)+v_{B}\mathbf{x}(B)+v_{C}\mathbf{x}(C)$, $\Phi$ must satisfy $2v_{A}>v_{B}+v_{C}$ and $2v_{B}>v_{A}$. In this case, $v_{A}=1,v_{B}=1,v_{C}=0$ works.
@@
The following is a variant of Farkas’ Lemma [farkas1902theorie], one of several similar “Theorems of the Alternative” stating that exactly one of two different linear systems has a solution. (See [mangasarian1994nonlinear, Chapter 2.4] for a list of such theorems.) A proof can be found in [gale1960theory, Theorem 2.10].
@@
We require the following discrete variant of Theorem˜6.3. The geometric intuition of this version is illustrated in Figure˜1.
Corollary 6.4.
cor:linear-alternative-integer Let $\mathbf{M}$ be a rational matrix. Exactly one of the following statements is true.
@@
Proof 6.5.
For convenience when we use LABEL:cor:linear-alternative-integer in proving LABEL:thm:potential-function-iff-entirely-execution-bounded, we swapped the roles of $\mathbf{u}$ and $\mathbf{v}$ in left- vs. right-multiplication with $\mathbf{M}$; the real-valued version of the statement of LABEL:cor:linear-alternative-integer is equivalent to Theorem˜6.3 by taking the transpose of $\mathbf{M}$.
@@
Figure 1: Geometric intuition of LABEL:cor:linear-alternative-integer. A matrix $\mathbf{M}$ has column vectors $\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3}$. The cone of $\mathbf{M}$ is the nonnegative span of these vectors, shown as a faded gray region. Exactly one of two scenarios occurs: a) The cone of $\mathbf{M}$ intersects the first quadrant (nonnegative orthant in higher dimensions) away from the origin, i.e., some semipositive point ($\mathbf{M}\mathbf{u}\geq\mathbf{0}$ above) is a nonnegative linear combination of vectors $\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3}$. b) The cone of $\mathbf{M}$ does not intersect the first quadrant except at the origin. In this case we can draw a dashed line (hyperplane in higher dimensions) separating the cone of $\mathbf{M}$ from the first quadrant. The orthogonal vector $\mathbf{v}\geq\mathbf{0}$ to this line lies in the first quadrant, but $\mathbf{v}\mathbf{M}<0$ means each vector $\mathbf{x}_{i}$ has negative dot product $\mathbf{v}\cdot\mathbf{x}_{i}<0$, i.e., all $\mathbf{x}_{i}$ vectors lie on the other side of the line.